OpenStudy (anonymous):

Plzhelpmethankyou John drove part of a 260 km trip at 80 km/h and the rest at 100 km/h. Find the distance he traveled at 80 km/h if his total driving time was 2 h and 30 min. I need help setting up the box/table like for rate x time = distance

OpenStudy (phi):

Here is the basic idea: rate x time = distance if you divide both sides by "rate" you get time = distance/rate you have two distances d1 (at 80 km/h) and d2 (at 100 km/h) you have two times: t1 = d1/80 and t2= d2/100 (using time= distance/rate) you know t1+t2= 2.5 hrs d1+d2= 260 km

OpenStudy (phi):

you can use d1+d2= 260 to write d2= 260-d1 (add -d1 to both sides) that makes t2= d2/100 = (260-d1)/100 (replace d2 with 260-d1 ) and t1 + t2 = 2.5 becomes d1/80 + (260-d1)/100 = 2.5 only 1 unknown value d1 is left. You can (if you know algebra) solve for d1

OpenStudy (phi):

I am not sure what you mean by setting up box/table, but I guess you mean how to set up the equations... and then how to solve them.